An economy has the production function y equals 20 times k super one solidus two full stop The current capital stock is 100, the depreciation rate is 10 percent, and the population growth rate is 2 percent. For income per person to grow, the saving rate must exceed

  1. 6 percent.
  2. 8 percent.
  3. 10 percent.
  4. 12 percent.
Answer #1

Answer

The correct answer is (a) 6 percent

Production function or Income is given by: y = 20k1/2.

We can see from above function, Output or Income will increase when k increases.

k will increase If Change in k >0 As, change in k = sy - (d + n)k

=> Change in k =  sy - (d + n)k > 0

where s = saving rate , d = depreciation rate = 10% = 0.10 , n = population growth rate = 2% = 0.02 and k = 100

Thus, sy - (d + n)k > 0

=> s(20k1/2)- (0.1 + 0.02)k > 0

=> s(20k1/2)- (0.1 + 0.02)k > 0

=> s(20*1001/2)- (0.1 + 0.02)100 > 0

=> 200s > 0.12*100

=> s > 12/200 = 0.06

=> saving rate > 6 percent.

Hence saving rate should be greater than 6%.

Hence, the correct answer is (a) 6 percent